# thanks: These two authors contributed equally to the workthanks: These two authors contributed equally to the work

### v1

@@ Abstract
Introduction – The fluorescence light emitted from a driven atom is characterised by periods of darkness interrupted by random, abrupt periods of brightness, corresponding to quantum jumps between atomic states [1, 2]. These early studies in atomic physics led to the first observations of quantum jumps [3, 4], which have become since then an experimental reality, observed in a variety of platforms in atomic physics, quantum optics and solid-state physics [5]. Quantum jumps also motivated the foundation of the modern theory of quantum trajectories [6, 7, 8], describing the physics of a quantum system which is continuously monitored by an external environment [9, 10].
@@
In recent years the effect of quantum jumps and quantum measurements have attracted fresh new interest in the context of open quantum many-body systems [11, 12] and their measurement-induced phase transitions [13, 14, 15]. Here the focus is on the entanglement properties of quantum many-body trajectories and how these are affected by the competition between unitary evolution and measurements [16].
@@
The standard setting where this transition has been studied involves the interplay between unitary evolution, described by a many-body Hamiltonian or by a random circuit, and a finite density of local projective measurements [17]. Alternatively, weak-measurement protocols have been also discussed, such as in the quantum jump or quantum state diffusion unravelling of the Lindblad evolution. In these cases one assume every site of the lattice to be exposed to a monitoring stochastic protocol. In all these examples the effect of measurements result in an entanglement transition from a volume-law scaling [18, 19, 20, 21, 22], or sub-volume in the case of non-interacting systems [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], to an area law where the system is frozen due to the quantum Zeno effect. This elusive criticality is hidden in the fluctuations of the monitoring protocol, while it remains transparent to the averaged state described by a Lindblad master equation. Experimental investigations of measurement-induced transitions have appeared [43, 44, 45] but are plagued by the postselection problem, i.e. the exponential overhead cost in sampling a quantum trajectory associated to a given measurement outcome, even though solutions in certain fine-tuned cases have been proposed [46, 47, 48].
@@
In this Letter, we explore the role of monitored dynamics and quantum jumps in a minimal model, where monitoring only occurs locally on a given site of an otherwise free fermionic chain. Local perturbations on quantum many-body systems, i.e. quantum impurities, have been long-known to induce non-trivial effects such as the ortoghonality catastrophe [49, 50, 51, 52]. Here we show that a non-unitary local monitoring results in a striking growth of entanglement entropy, which slowly saturates to an extensive value scaling with the volume of the subsystem size. This is in stark contrast with both the unitary case and the fully postselected, no-click limit. Therefore our results show that quantum jumps are crucial in generating entanglement. We explain this result by looking at the statistics of quantum jumps, which we reveal to display periods of bunching, interspersed with long dark intervals, similar to the resonance fluorescence of a driven atom. We demonstrate that the presence of these dark periods is the key to extensive entanglement generation and that by monitoring more sites the entanglement generally decreases, recovering the area-law regime expected when each lattice site is monitored.
@@
Model and Monitoring Protocol – We consider a non-interacting $1D$ fermionic system coupled to a single monitoring device at site $j_{0}$. The fermionic model can be described by the Hamiltonian
@@
with $L$ sites and periodic boundary conditions, $J$ being the hopping strength. We consider a monitoring quantum jump protocol where the system wave function evolves randomly in time according to the stochastic Schrödinger equation [7, 6, 8, 10]
$\partial_{t}\ket{\psi}$ $=-idt\left\{H-\frac{i}{2}\left(L^{\dagger}_{j_{0}}L_{j_{0}}-\langle L^{\dagger}_{j_{0}}L_{j_{0}}\rangle\right)\right\}\ket{\psi}$
@@
where $d\xi(t)\in\{0,1\}$ is an inhomogeneous Poisson increment such that $P(d\xi(t)=1)=dt\bra{\psi(t)}L_{j_{0}}^{\dagger}L_{j_{0}}\ket{\psi(t)}$. The action of the measurement device can be described by one jump operator of the form $L_{j_{0}}=\sqrt{\gamma}n_{j_{0}}$ (as depicted in the sketch of Fig. 1 (a) ). The first line of Eq. (2) describes the deterministic evolution where unitary dynamics is modified by the measurement back-action, leading to a non-Hermitian Hamiltonian of the form
$H_{NH}=-J\sum_{j=0}^{L-1}c_{j+1}^{\dagger}c_{j}+c_{j}^{\dagger}c_{j+1}-i\frac{\gamma}{2}n_{j_{0}}.$ (3)
This Hamiltonian drives the dynamics in the no-click limit, corresponding to post-selecting the trajectory where no quantum jump arises during the evolution, or in between consecutive quantum jumps. Eq. (3) describes free fermions in an imaginary scattering potential, a problem which has been previously studied [53, 54]. A generic quantum trajectory results in a series of quenched evolution under the effective non-Hermitian Hamiltonian. For each realisation of this stochastic process, we obtain a trajectory that can be numerically efficiently computed for Gaussian systems [16]. On the other hand, averaging over the full set of trajectories yields a Lindblad master equation dynamics for a free fermion chain subject to a local dephasing process at site $j_{0}$; a problem addressed in Ref. [54]. In this work we are interested in studying the entanglement structure of quantum trajectories evolving under the protocol described above. Since the evolution conserves the purity and the initial state is chosen to be pure we can consider the von Neumann entanglement entropy defined as $S_{\xi}(t)=-\mathrm{Tr}_{A}\left[\rho_{\xi}^{A}(t)\ln\rho_{\xi}^{A}(t)\right]$ where we have introduced a bi-partition A,B (cf Fig. 1 a) ), with the reduced density matrix $\rho^{A}_{\xi}(t)=\mathrm{Tr}_{B}\lvert\psi_{\xi}(t)\rangle\langle\psi_{\xi}(t)\rvert$. In the following, we will discuss in particular the behavior of the entanglement entropy averaged over the measurement noise as
@@
We choose, as an initial state, a Fermi sea with constant filling, $\nu$. This choice makes the no-click evolution a local quench, since it is an eigenstate of the Hermitian part of $H_{NH}$.
Entanglement properties – We start by discussing the dynamics of entanglement entropy, plotted in Fig. 1c for $\gamma=0.5$ and different subsystem sizes $\ell$. In order to highlight the growth due to the monitoring we subtract the value of the entropy at time $t=0$, corresponding to the critical Fermi sea. We begin by observing a remarkable production of entanglement and a threshold behavior. After some initial transient regime, where the entanglement entropy remains close to the initial value, we see a sizeable growth with time and a strong dependence on subsystem size. Furthermore, the entanglement entropy reaches a steady state on a notably slow timescale. While this timescale for free fermions is usually proportional to the sub-system length, here we find it to be much longer and compatible with an exponential scaling [55]. Another remarkable feature is that the steady-state value of the entanglement entropy grows linearly with the subsystem size $\ell$, as illustrated in Fig. 1c, i.e. it displays volume law scaling. Unlike typical free-fermion MIPT [23, 32, 25], where the entire chain is monitored, we do not observe an area law scaling regime, even at reasonably large measurement rates $\gamma$. Crucially, this is a feature of the local monitoring protocol: indeed, if we increase the number of sites which are monitored we see that the entanglement entropy decreases and the volume law scaling turns into an area-law one [55]. The dramatic effect of a single monitoring protocol on the entanglement dynamics is one of the major results of this work. These results are particularly striking when compared to the no-click limit evolution, corresponding to a local quench by an imaginary scattering potential, but also if compared to the analogous unitary dynamics, i.e. when the quench is realized by a real scattering potential. In both cases indeed the entanglement entropy would quickly saturate to a value independent on sub-system size [55].
@@
Figure 2: Quantum trajectories at fixed waiting time. (a) Entanglement dynamics of quantum trajectories trajectories where the waiting time is fixed to $\Delta T$. Dark curves represent random typical trajectories and the inset shows a zoom on the dynamics of different waiting times $\Delta T$. (b) Rate of Entanglement growth $v(\Delta T)$ of quantum trajectories as a function of the imposed waiting time. The vertical dashed line indicates the average waiting time $\overline{\tau}$ of the long tail of the full WTD. Parameters: $\gamma=0.5,\,J=0.5,\,j_{0}=0,\,L=128,\,\nu=1/4$.
Figure 3: Connection between the tight-binding energy and the entanglement growth. (a) Dynamics of the tight binding energy $E(t)$ (red curve) and of the entanglement entropy $S(t)$ (black curve). The vertical dashed gray lines denote quantum jumps events. (b) Analysis of the correlations between the change of entanglement $\Delta S$ and the variation of energy $\Delta E_{\mathrm{qj}}$ The top-right panel is a sketch defining the different quantities, thereby $\Delta S_{\mathrm{qj,n}}$ denotes the change of entanglement during the $n$-th quantum jump, and $\tau_{qj,n}$ is the waiting time following the $n$-th jump. The bottom-left panel shows the correlation between the variation of energy during a quantum jump $\Delta E_{\mathrm{qj}}$ and the change of entanglement $\Delta S=\Delta S_{\mathrm{qj}}+\Delta S_{\mathrm{nH}}$, happening between the same jumps and the following one (in this expression $\Delta S_{\mathrm{nH}}$ is the contribution to the entanglement produced in the non-Hermitian evolution between these two jumps, and $n$ is omitted whenever all quantities refer to the same jump event). The dark line represents the linear trend. In top-left panel we show the marginal distribution $P(\Delta S)$ conditioned on the energy variation being larger or smaller than a threshold (dotted line of bottom-left panel). The bottom-right panel shows the marginal distribution $P(\Delta E_{\mathrm{qj}})$. (c) Waiting time in function of the variation of energy $\Delta E_{\mathrm{qj},n}$ occurring in the jump following the wait $\tau_{\mathrm{qj},n-1}$. Parameters: $\gamma=0.5,\,J=0.5,\,j_{0}=0,\,L=128,\,\nu=1/4$.
@@
Statistics of Jumps Waiting times – Looking at a characteristic single shot evolution of the entanglement entropy, plotted in Fig. 1b, already suggests something peculiar occurs in this problem. The way in which quantum jumps are distributed displays two alternating behaviors, characterized by bunching periods and dark periods, respectively. Specifically, the dynamics appears to consist of series of quantum jumps (plotted as orange dots) occurring in close succession (bunching periods), interspersed with long intervals during which the system does not jump, and the dynamics is governed by non-Hermitian evolution (dark periods). This intuition is confirmed by the waiting time distribution (WTD) of quantum jumps, displayed in Fig. 1d, which exhibits two distinct timescales, thereby showing a significant deviation from the typical Poissonian distribution seen in jumps unraveling  [^1] The statistics of jumps do not always follow a Poissonian distribution throughout the entire dynamics; however, it usually is reached in the steady state. This is not the case here. Most quantum jumps occur at very short waiting time $\tau$, following a Poissonian distribution. However, we also observe a sizable amount of quantum jumps occurring at a much longer waiting times, giving rise to a statistically significant tails in the distribution. This kind of distribution is known as super-Poissonian, meaning that the ratio of the variance $\sigma$ to the mean $\mu$ is greater than one (where a Poissonian distribution has a ratio of one). Super-Poissonian distributions often arise in contexts where events are bunched. We will explain how this distinctive feature plays a key role in generating the observed entanglement dynamics.
@@
Entanglement growth, dark periods and bath disentangling – The emergence of these two time scales in the WTD enables the presence of dark intervals in the monitoring dynamics. The jumps that occur during the bunching periods contribute to the initial exponential decay in the WTD (see  Fig.1 d), while the jumps following the dark periods are responsible for the long tail. An interesting aspect to discern is whether the increase in entanglement can be associated either to the presence of bunching or to the long periods during which the measurement device does not click.
@@
We now argue that the key to the extensive volume-law like entanglement growth in our problem of local monitoring consist in these dark periods during which the evolution is driven by a non-Hermitian Hamiltonian, starting however from a different initial state after each random quantum jump. To investigate this, in Fig. 2 we analyze trajectories with fixed waiting times $\Delta T$  [^2] We generate trajectories with fixed jumping times, irrespectively of each jumping probability. In principle this same collection of trajectories could have been obtained through post-selection, but it would have been numerically impractical. We stress that the two procedures converge to the same result., and we evaluate the entanglement dynamics along them. As we show in panel b, the rate of entanglement growth, $v(\Delta T)$, displays a clear non-monotonic dependence on the fixed waiting time $\Delta T$. Importantly, we see that $v(\Delta T)$ reaches its optimal value when $\Delta T$ matches the average waiting time $\overline{\tau}$ observed during the dark intervals (cf. Fig 1d). This analysis shows that the essential factor for achieving significant entanglement growth is the presence of these dark periods. To gain further insight into why long dark periods enable entanglement growth we now look in more details at the statistics of the entanglement entropy changes along a quantum trajectory. This metric was recently introduced to understand the mutual role of non-Hermitian evolution and quantum jumps for entanglement phase transitions [16]. As we show in  [55], when applied to our model this analysis reveals that the net effect of quantum jump events is to decrease the entanglement entropy, while the non-Hermitian evolution is responsible for its growth. Interestingly the entanglement loss due to a jump depends non-monotonously on the length of the waiting time preceding the event [55], suggesting that bunched jumps have less effect than the one happening after a dark period of order $\overline{\tau}$.
@@
We can now understand at least qualitatively the origin of this extensive entanglement production: when a quantum jump hits the monitored site after a long dark period of no-clicks it induces a non-trivial effect across the chain, disentangling part of it and increasing its energy, thus bringing the system far away from equilibrium. As a result, the non-Hermitian evolution which follows this type of events can be considered akin to a global quantum quench, where part of the chain is perturbed and re-entangled rapidly and efficiently, rather than a local one where entanglement entropy growth can be, at best, logarithmic. Out of this balance between loss due to quantum jumps and gain due to entangling evolution, a steady state is eventually reached after a certain number of dark periods. In the no-click evolution, on the other hand, although the dynamics is still driven by the same non-Hermitian Hamiltonian, the initial state is an eigenstate of the free-fermion Hamiltonian. This implies that the perturbation to the state is only due to the measurement back-action (responsible for the non-Hermiticity), whose effect is too local to induce any sizable change in the bulk of the chain. As such it is not surprising that the entanglement production in the no-click limit remains small [55]. This also clarifies why the entanglement production is non-monotonous with $\Delta T$: if we increase $\Delta T$ too much, the system remains stuck with the non-Hermitian Hamiltonian in this steady state for too long, slowing down the entanglement production.Our results therefore showcase how local measurements can have dramatic effect on many-body states. Importantly, while our discussion has been focused on the quantum jump protocol the main effect survives also in presence of projective measurements [55].
@@
Conclusions – In this work we have discussed the effect of local on-site monitoring via quantum jumps on a fermionic chain. Remarkably, we have shown that such a local non-unitary perturbation results in a slow growth of entanglement entropy which saturates to a volume law value, in striking contrast to both the unitary case and the no-click limit where only subextensive entanglement can be generated by local perturbations. We have highlighted the key role of quantum jumps and their super-Poissonian WTD in explaining the entanglement production. In particular we have shown that, while quantum jumps decrease the entanglement entropy they also reset the many-body state disentangling the monitored site and its surrounding from the rest of the chain and thus leaving space for the non-Hermitian evolution to produce entanglement during the long dark intervals. Our results represent one of the simplest yet nontrivial effects due to monitoring, which crucially comes with low overhead in postselection, given the local monitoring and the slow dynamics leading to few quantum jumps. As such, our results could be experimentally verified in quantum simulators with on-site addressing, such as Rydberg arrays [58, 59, 60].

### v2

@@ Abstract
Introduction – The fluorescence light emitted from a driven atom is characterised by periods of darkness interrupted by random, abrupt periods of brightness, corresponding to quantum jumps between atomic states Cohen-Tannoudji and Dalibard (1986); Cook and Kimble (1985). These early studies in atomic physics led to the first observations of quantum jumps Bergquist et al. (1986); Nagourney et al. (1986), which have become since then an experimental reality, observed in a variety of platforms in atomic physics, quantum optics and solid-state physics Minev et al. (2019). Quantum jumps also motivated the foundation of the modern theory of quantum trajectories Ueda (1990); Dalibard et al. (1992); Gardiner et al. (1992), describing the physics of a quantum system which is continuously monitored by an external environment Plenio and Knight (1998); Wiseman and Milburn (2009).
@@
In recent years the effect of quantum jumps and quantum measurements have attracted fresh new interest in the context of open quantum many-body systems Daley (2014); Fazio et al. (2024) and their measurement-induced phase transitions Nahum et al. (2021); Li et al. (2018, 2019). Here the focus is on the entanglement properties of quantum many-body trajectories and how these are affected by the competition between unitary evolution and measurements Le Gal et al. (2024).
@@
The standard setting where this transition has been studied involves the interplay between unitary evolution, described by a many-body Hamiltonian or by a random circuit, and a finite density of local projective measurements Fisher et al. (2023). Alternatively, weak-measurement protocols have been also discussed, such as in the quantum jump or quantum state diffusion unravelling of the Lindblad evolution. In these cases one assume every site of the lattice to be exposed to a monitoring stochastic protocol. In all these examples the effect of measurements result in an entanglement transition from a volume-law scaling Fuji and Ashida (2020); Lunt and Pal (2020); Doggen et al. (2022); Xing et al. (2024); Altland et al. (2022), or sub-volume in the case of non-interacting systems Cao et al. (2019); Fidkowski et al. (2021); Coppola et al. (2022); Lóio et al. (2023); Poboiko et al. (2023); Jian et al. (2023); Fava et al. (2023); Carisch et al. (2023); Jin and Martin (2023); Alberton et al. (2021); Van Regemortel et al. (2021); Turkeshi et al. (2021); Botzung et al. (2021); Bao et al. (2021); Turkeshi et al. (2022); Piccitto et al. (2022); Kells et al. (2023); Paviglianiti and Silva (2023); Müller et al. (2022); Soares et al. (2024), to an area law where the system is frozen due to the quantum Zeno effect. This elusive criticality is hidden in the fluctuations of the monitoring protocol, while it remains transparent to the averaged state described by a Lindblad master equation. Experimental investigations of measurement-induced transitions have appeared Noel et al. (2022); Koh et al. (2023); Google AI and Collaborators (2023) but are plagued by the postselection problem, i.e. the exponential overhead cost in sampling a quantum trajectory associated to a given measurement outcome, even though solutions in certain fine-tuned cases have been proposed Ippoliti and Khemani (2021); Passarelli et al. (2024); Garratt and Altman (2024).
@@
In this Letter, we explore the role of monitored dynamics and quantum jumps in a minimal model, where monitoring only occurs locally on a given site of an otherwise free fermionic chain. Local perturbations on quantum many-body systems, i.e. quantum impurities, have been long-known to induce non-trivial effects such as the ortoghonality catastrophe Anderson (1967); Bettelheim et al. (2006); Schiró and Mitra (2014); Tonielli et al. (2019). Here we show that a non-unitary local monitoring results in a striking growth of entanglement entropy, which slowly saturates to an extensive value scaling with the volume of the subsystem size. This is in stark contrast with both the unitary case and the fully postselected, no-click limit. Therefore our results show that quantum jumps are crucial in generating entanglement. We explain this result by looking at the statistics of quantum jumps, which we reveal to display periods of bunching, interspersed with long dark intervals, similar to the resonance fluorescence of a driven atom. We demonstrate that the presence of these dark periods is the key to extensive entanglement generation and that by monitoring more sites the entanglement generally decreases, recovering the area-law regime expected when each lattice site is monitored.
@@
Model and Monitoring Protocol – We consider a non-interacting $1D$ fermionic system coupled to a single monitoring device at site $j_{0}$. The model can be described by the Hamiltonian
@@
with $L$ sites and periodic boundary conditions, $J$ being the hopping strength. We consider a monitoring quantum jump protocol where the system wave function evolves randomly in time according to the stochastic Schrödinger equation Dalibard et al. (1992); Ueda (1990); Gardiner et al. (1992); Wiseman and Milburn (2009)
$\partial_{t}\ket{\psi}$ $=-idt\left\{H_{0}-\frac{i}{2}\left(L^{\dagger}_{j_{0}}L_{j_{0}}-\langle L^{\dagger}_{j_{0}}L_{j_{0}}\rangle\right)\right\}\ket{\psi}$
@@
where $d\xi(t)\in\{0,1\}$ is an inhomogeneous Poisson increment such that $P(d\xi(t)=1)=dt\bra{\psi(t)}L_{j_{0}}^{\dagger}L_{j_{0}}\ket{\psi(t)}$. The action of the measurement device can be described by one jump operator of the form $L_{j_{0}}=\sqrt{\gamma}n_{j_{0}}$ (as depicted in the sketch of Fig. 1 (a) ).
A generic quantum trajectory described in Eq. (2) results in a series of evolutions under an effective non-Hermitian Hamiltonian $H_{NH}=H_{0}-i\frac{\gamma}{2}n_{j_{0}}$, interrupted by random quantum jumps where the number of fermions at site $j_{0}$ is measured. Two limits of this problem have been studied: (i) the fully post-selected evolution corresponding to a trajectory with no quantum jumps, also called the no-click limit, where the problem reduces to free fermions in an imaginary scattering potential Stefanini and Marino (2023); Dolgirev et al. (2020) and (ii) the dynamics of the average state described by a Lindblad master equation for a free fermion chain subject to a local dephasing Dolgirev et al. (2020). In this work we are interested in studying the entanglement structure of quantum trajectories evolving under the protocol described above. Since the evolution conserves the purity and the initial state is chosen to be pure we can consider the von Neumann entanglement entropy defined as $S_{\xi}(t)=-\mathrm{Tr}_{A}\left[\rho_{\xi}^{A}(t)\ln\rho_{\xi}^{A}(t)\right]$ where we have introduced a bi-partition A,B (cf Fig. 1 (a) ), with the reduced density matrix $\rho^{A}_{\xi}(t)=\mathrm{Tr}_{B}\lvert\psi_{\xi}(t)\rangle\langle\psi_{\xi}(t)\rvert$. In the following, we will discuss in particular the behavior of the entanglement entropy averaged over the measurement noise as
@@
We choose, as an initial state, a Fermi sea with constant filling, $\nu$. This choice makes the no-click evolution a local quench, since it is an eigenstate of the Hermitian part of $H_{NH}$. Since the system remains Gaussian along the quantum trajectory, Eq. (2) can be efficiently simulated Le Gal et al. (2024).
Entanglement properties – We start by discussing the dynamics of entanglement entropy, plotted in Fig. 1(c) for $\gamma=0.5$ and different subsystem sizes $\ell$. In order to highlight the growth due to the monitoring we subtract the value of the entropy at time $t=0$, corresponding to the critical Fermi sea. We begin by observing a remarkable production of entanglement and a threshold behavior. After some initial transient regime, where the entanglement entropy remains close to the initial value, we see a sizeable growth with time and a strong dependence on subsystem size. Furthermore, the entanglement entropy reaches a steady state on a notably slow timescale. While this timescale for free fermions is usually proportional to the sub-system length, here we find it to be much longer and compatible with an exponential scaling  [^1] Supplemental Materials. Another remarkable feature is that the steady-state value of the entanglement entropy grows linearly with the subsystem size $\ell$, as illustrated in Fig. 1 (c), i.e. it displays volume law scaling. Unlike typical free-fermion MIPT Cao et al. (2019); Alberton et al. (2021); Coppola et al. (2022), where the entire chain is monitored, we do not observe an area law scaling regime, even at reasonably large measurement rates $\gamma$. Crucially, this is a feature of the local monitoring protocol: indeed, if we increase the number of sites which are monitored we see that the entanglement entropy decreases and the volume law scaling turns into an area-law one Note (1). The dramatic effect of a single monitoring protocol on the entanglement dynamics is one of the major results of this work. These results are particularly striking when compared to the no-click limit evolution, corresponding to a local quench by an imaginary scattering potential, but also if compared to the analogous unitary dynamics, i.e. when the quench is realized by a real scattering potential. In both cases indeed the entanglement entropy would quickly saturate to a value independent on sub-system size Note (1).
@@
Statistics of Jumps Waiting times – Looking at a characteristic single shot evolution of the entanglement entropy, plotted in Fig. 1 (b), already suggests something peculiar occurs in this problem. The way in which quantum jumps are distributed displays two alternating behaviors, characterized by bunching periods and dark periods, respectively. Specifically, the dynamics appears to consist of series of quantum jumps (plotted as orange dots) occurring in close succession (bunching periods), interspersed with long intervals during which the system does not jump, and the dynamics is governed by non-Hermitian evolution (dark periods). This intuition is confirmed by the waiting time distribution (WTD) of quantum jumps, displayed in Fig. 1 (d), which exhibits two distinct timescales, thereby showing a significant deviation from the typical Poissonian distribution seen in jumps unraveling  [^2] The statistics of jumps do not always follow a Poissonian distribution throughout the entire dynamics; however, it usually is reached in the steady state. This is not the case here. Most quantum jumps occur at very short waiting time $\tau$, following a Poissonian distribution. However, we also observe a sizable amount of quantum jumps occurring at a much longer waiting times, giving rise to a statistically significant tails in the distribution. This kind of distribution is known as super-Poissonian, meaning that the ratio of the variance $\sigma$ to the mean $\mu$ is greater than one (where a Poissonian distribution has a ratio of one). Super-Poissonian distributions often arise in contexts where events are bunched. We will explain how this distinctive feature plays a key role in generating the observed entanglement dynamics.
@@
Entanglement growth, dark periods and bath disentangling – The emergence of these two time scales in the WTD enables the presence of dark intervals in the monitoring dynamics. The jumps that occur during the bunching periods contribute to the initial exponential decay in the WTD (see  Fig.1 (d)), while the jumps following the dark periods are responsible for the long tail. An interesting aspect to discern is whether the increase in entanglement can be associated either to the presence of bunching or to the long periods during which the measurement device does not click.
Figure 2: Entanglement growth mechanism. (a) Entanglement dynamics of quantum trajectories trajectories where the waiting time is fixed to $\Delta T$. Dark curves are typical random trajectories and the inset shows rate of entanglement growth $v(\Delta T)$ where $\overline{\tau}$ is the average waiting time of the long tail of the full WTD. (b) Dynamics of the tight binding energy $E(t)$ (red curve) and of the entanglement entropy $S(t)$ (black curve). (c) Sketch defining the different quantities, thereby $\Delta S_{\mathrm{qj}}$ denotes the change of entanglement during one QJ, and $\tau_{qj}$ is the waiting time following a jump. (d) Marginal distribution $P(\Delta S)$ of the change of entanglement $\Delta S=\Delta S_{\mathrm{qj}}+\Delta S_{\mathrm{nH}}$ conditioned on the energy variation being larger or smaller $\Delta E_{\mathrm{cut}}$ (e) Distribution $P(\Delta E_{\mathrm{qj}})$ of the change of energy during a QJ. Parameters: $\gamma=0.5,\,J=0.5,\,j_{0}=0,\,L=128,\,\nu=1/4$.
@@
We now argue that the key to the extensive volume-law like entanglement growth in our problem of local monitoring consist in these dark periods during which the evolution is driven by a non-Hermitian Hamiltonian, starting however from a different initial state after each random quantum jump. To investigate this, in Fig. 2 (a) we analyze trajectories with fixed waiting times $\Delta T$  [^3] We generate trajectories with fixed jumping times, irrespectively of each jumping probability. In principle this same collection of trajectories could have been obtained through post-selection, but it would have been numerically impractical. We stress that the two procedures converge to the same result., and we evaluate the entanglement dynamics along them. As we show in the inset, the rate of entanglement growth, $v(\Delta T)$, displays a clear non-monotonic dependence on the fixed waiting time $\Delta T$. Importantly, we see that $v(\Delta T)$ reaches its optimal value when $\Delta T$ matches the average waiting time $\overline{\tau}$ observed during the dark intervals (cf. Fig. 1 (d)). This analysis shows that the essential factor for achieving significant entanglement growth is the presence of these dark periods. To gain further insight into why long dark periods enable entanglement growth we now look in more details at the statistics of the entanglement entropy changes along a quantum trajectory. This metric was recently introduced to understand the mutual role of non-Hermitian evolution and quantum jumps for entanglement phase transitions Le Gal et al. (2024). As we show in  Note (1), when applied to our model this analysis reveals that the net effect of quantum jump events is to decrease the entanglement entropy, while the non-Hermitian evolution is responsible for its growth. Interestingly the entanglement loss due to a jump depends non-monotonously on the length of the waiting time preceding the event Note (1), suggesting that bunched jumps have less effect than the one happening after a dark period of order $\overline{\tau}$.
@@
We can now understand at least qualitatively the origin of this extensive entanglement production: when a quantum jump hits the monitored site after a long dark period of no-clicks it induces a non-trivial effect across the chain, disentangling part of it and increasing its energy, thus bringing the system far away from equilibrium. As a result, the non-Hermitian evolution which follows this type of events can be considered akin to a global quantum quench, where part of the chain is perturbed and re-entangled rapidly and efficiently, rather than a local one where entanglement entropy growth can be, at best, logarithmic. Out of this balance between loss due to quantum jumps and gain due to entangling evolution, a steady state is eventually reached after a certain number of dark periods. In the no-click evolution, on the other hand, although the dynamics is still driven by the same non-Hermitian Hamiltonian, the initial state is an eigenstate of the free-fermion Hamiltonian. This implies that the perturbation to the state is only due to the measurement back-action (responsible for the non-Hermiticity), whose effect is too local to induce any sizable change in the bulk of the chain. As such it is not surprising that the entanglement production in the no-click limit remains small Note (1). This also clarifies why the entanglement production is non-monotonous with $\Delta T$: if we increase $\Delta T$ too much, the system remains stuck with the non-Hermitian Hamiltonian in this steady state for too long, slowing down the entanglement production.Our results therefore showcase how local measurements can have dramatic effect on many-body states. Importantly, while our discussion has been focused on the quantum jump protocol the main effect survives also in presence of projective measurements Note (1).
@@
Conclusions – In this work we have discussed the effect of local on-site monitoring via quantum jumps on a fermionic chain. Remarkably, we have shown that such a local non-unitary perturbation results in a slow growth of entanglement entropy which saturates to a volume law value, in striking contrast to both the unitary case and the no-click limit where only subextensive entanglement can be generated by local perturbations. We have highlighted the key role of quantum jumps and their super-Poissonian WTD in explaining the entanglement production. In particular we have shown that, while quantum jumps decrease the entanglement entropy they also reset the many-body state disentangling the monitored site and its surrounding from the rest of the chain and thus leaving space for the non-Hermitian evolution to produce entanglement during the long dark intervals. Our results represent one of the simplest yet nontrivial effects due to monitoring, which crucially comes with low overhead in postselection, given the local monitoring and the slow dynamics leading to few quantum jumps. As such, our results could be experimentally verified in quantum simulators with on-site addressing, such as Rydberg arrays Nguyen et al. (2018); Ravon et al. (2023); Browaeys and Lahaye (2020).
Acknowledgements.